Tanner is three times as old as Rachel. The sum of their age is 160. How old are each of them?

Answer by ikleyn(52788)

.
Tanner + Rashel = 160.


3R      + R     = 160


4R = 160  ====>  R = 160/4 = 40.


Answer.  R is 40 years old.   T is 3*40 = 120 years old.